RS1 remaining parent material is 12.5%. RS2 remaining parent material is 25%. Based on the amount of remaining parent material in the rocks, how many half-lives have occurred for each?
1
Expert's answer
2019-04-15T13:36:45-0400

"RS1:\;\;(\frac{1}{2})^n=0.125=\frac{1}{8} \;\; ->\;\; n=3.\;\;3\;half-lives\;occured.\\
RS2:\;\;(\frac{1}{2})^n=0.25=\frac{1}{4} \;\; ->\;\; n=2.\;\;2\;half-lives\;occured."